# Enhanced simultaneous quantum-classical communications under composable security

### v1

@@ B.2 Renormalisation
We are given some freedom in the choice of our normalisation. Recalling that $c_{d}=(1-\delta)c$ for $10\leq\delta<$ 1, one potential option is defining
@@
$\bm{\mu}_{d}^{\prime}$ $=\left(0,0,\frac{2\sqrt{T}de_{C}}{\sqrt{2(1-\delta)^{2}}},\frac{2\sqrt{T}de_{C}}{\sqrt{2(1-\delta)^{2}}}\right)^{T},$ (82)
$\bm{V}_{d}^{\prime}$ $=\begin{pmatrix}a\mathbb{I}&c\sigma_{Z}\\ c\sigma_{Z}&\left[\frac{b_{d}+1}{(1-\delta)^{2}}-1\right]\mathbb{I}\end{pmatrix}.$ (83)
@@
However, the specific choice of renormalisation $\Delta_{V}\leavevmode\nobreak\ =\leavevmode\nobreak\ (1-\delta)^{2}$, such that $c^{\prime}=c$, is by no means the optimal choice. An alternative solution might be to choose $\Delta_{V}$ such that Bob’s covariance matrix element is invariant under the postprocessing and rescaling, i.e.
@@ Appendix C Asymptotic key rate calculation
$\bm{V}_{AB}$ $=\begin{pmatrix}a\mathbb{I}&c\sigma_{Z}\\ c\sigma_{Z}&b\mathbb{I}\end{pmatrix},$ (98)
then the von Neumann entropies $S(\hat{\rho}_{G}^{AB})$ and $S(\hat{\rho}_{G}^{A|B})$ may be calculated from the symplectic eigenvalues $\lambda_{1,2}$ and $\lambda_{3}$ of the covariance matrices $\bm{V}_{AB}$ and $\bm{V}_{A|B}$ respectively [22]:
@@ Appendix D Parameter estimation and security parameters in the finite-size protocol
$\bm{V}_{AB}$ $=\begin{pmatrix}a\mathbb{I}&c\sigma_{Z}\\ c\sigma_{Z}&b\mathbb{I}\end{pmatrix},$ (110)
@@ Appendix E Satellite-based platforms
[…]

### v2

@@ B.2 Renormalisation
We are given some freedom in the choice of our normalisation. Recalling that $c_{d}=(1-\delta)c$ for $0\leq\delta<1$, one potential option is defining
@@
$\boldsymbol{\mu}_{d}^{\prime}$ $=\left(0,0,\frac{2\sqrt{T}de_{C}}{\sqrt{2(1-\delta)^{2}}},\frac{2\sqrt{T}de_{C}}{\sqrt{2(1-\delta)^{2}}}\right)^{T},$ (82)
$\boldsymbol{V}_{d}^{\prime}$ $=\begin{pmatrix}a\mathbb{I}&c\sigma_{Z}\\ c\sigma_{Z}&\left[\frac{b_{d}+1}{(1-\delta)^{2}}-1\right]\mathbb{I}\end{pmatrix}$. (83)
@@
However, the specific choice of renormalisation $\Delta_{V}\penalty 10000\ =\penalty 10000\ (1-\delta)^{2}$, such that $c^{\prime}=c$, is by no means the optimal choice. An alternative solution might be to choose $\Delta_{V}$ such that Bob’s covariance matrix element is invariant under the postprocessing and rescaling, i.e.
@@ Appendix C Asymptotic key rate calculation
$\boldsymbol{V}_{AB}$ $=\begin{pmatrix}a\mathbb{I}&c\sigma_{Z}\\ c\sigma_{Z}&b\mathbb{I}\end{pmatrix}$, (98)
then the von Neumann entropies $S(\hat{\rho}_{G}^{AB})$ and $S(\hat{\rho}_{G}^{A|B})$ may be calculated from the symplectic eigenvalues $\lambda_{1,2}$ and $\lambda_{3}$ of the covariance matrices $\boldsymbol{V}_{AB}$ and $\boldsymbol{V}_{A|B}$ respectively [22]:
@@ Appendix D Parameter estimation and security parameters in the finite-size protocol
$\boldsymbol{V}_{AB}$ $=\begin{pmatrix}a\mathbb{I}&c\sigma_{Z}\\ c\sigma_{Z}&b\mathbb{I}\end{pmatrix}$, (110)
@@ Appendix E Satellite-based platforms
However, the update of channel information is limited by the satellite-to-Earth round-trip time, which is ${\sim}3.3\,\text{ms}$, assuming that a typical low-Earth-orbit satellite (positioned at an altitude of $500\,\text{km}$) is located at the zenith point relative to the ground station. Since the coherence time of an atmospheric channel is typically on the order of $1\,\text{ms}$ [33], such an implementation can be very challenging. Nevertheless, while the requirement of precise information on the instantaneous channel transmissivity is strictly necessary at Bob, it is expected that this requirement can be relaxed at Alice in a downlink scenario (where channel fluctuations are minimal). Therefore, SQCC should be suitable for the downlink configuration, which represents the mainstream configuration adopted in satellite-based optical classical and quantum communications. It should also be noted that SQCC is naturally well-suited for the inter-satellite configuration (an important yet less explored scenario for communications within a satellite constellation) due to the absence of atmospheric turbulence in space. To summarize, within the context of satellite-based communications, the results presented in this work will be useful for both satellite-to-Earth (downlink) and inter-satellite configurations. They will not be useful for the Earth-to-satellite (uplink) configuration.